Determine the cumulative distribution function for following:(2 Marks) F(x) 0 x < 0 0.2 x 0 <= x < 4 0.04 x + 0.64 4 <= x < 9 1 9 <= x


1
Expert's answer
2022-06-13T15:23:44-0400

The cumulative distribution function


"F(x)= \begin{cases}
 0 &x<0 \\
 0.2x &0\le x<4\\
 0.04x+0.64 & 4\le x<9\\
 1& x\ge9\\
\end{cases}"

Determine the probability density function


"f(x)= \begin{cases}
 0 &x<0 \\
 0.2 &0\le x<4\\
 0.04 & 4\le x<9\\
\end{cases}"